lim e7s+2

s→–œ



1
Expert's answer
2021-02-24T06:02:46-0500

Ans:-

"lim_{s\rightarrow -\infty}e^{7s}+2" "\Rightarrow" "e^{-\infty} +2" "\Rightarrow" 2 ( since "e^{-\infty}= 0" )